OpenStudy (anonymous):

@xGuardians @Hero One leg in a right triangle is 9 inches long. The other leg is 12 inches long. How long is the hypotenuse? A. 15 inches B. 16 inches C. 18 inches D. 21 inches <--- The biggest square should have the exact same area as the other two squares put together. 12+9=21

hero (hero):

Face palm

OpenStudy (anonymous):

hmm no you need to solve \[9^2+12^2=h^2\]

hero (hero):

9 and 12 are legs, not areas

OpenStudy (anonymous):

I got is from here. http://www.mathsisfun.com/pythagoras.html I don't see what I did wrong. @xGuardians

OpenStudy (xguardians):

9² + 12² = C² (Hypotenuse)

OpenStudy (xguardians):

You get that right?

OpenStudy (anonymous):

OK. :)

OpenStudy (xguardians):

81 + 144 = C²

OpenStudy (xguardians):

You can't just square root them all right away, that's your mistake.

OpenStudy (xguardians):

225 = C² C = 15

OpenStudy (xguardians):

You have to add the two together before you can square root, hope this helps!

OpenStudy (anonymous):

So... I think I get it... 9^2 = 81 12^2 = 144 81+144=225 15^2 = 225 I'm positive that's right this time. :) @xGuardians

OpenStudy (anonymous):

now you get the square root of 225

OpenStudy (xguardians):

Yes so the answer is A. 15 :)

OpenStudy (anonymous):

Thanks!